a uniformly acceletated body travels during the first two intervals of 6s distances of 30m and 80m.Find the distance travelled by the body at 20th second

Grade:11

1 Answers

Arun
5 years ago
 
Dear student
 
s = ut + ½ at^2
30 = 6u + 18a
 
Also
110 = 12u + 72 a
 
on solving
36 a = 50
a = 50/36 = 25/18
 
also u= 5/6
 
hence distance travelled i. 20th second = u + ½ a (2 t – 1)
 
 = 5/6 + ½ * 18/25 ( 39)
 = 0.833 + (14.04)
 
 = 14.873